OpenStudy (girlstudy):

A ball's mass is 1.50 kilograms, and it leaves the racket with a horizontal speed of, say, 90 mph. Find the average force on the tennis ball as it is served.

OpenStudy (mhchen):

I'm assuming you the 'horizontal speed' is only the x-component of the actual force right'?

OpenStudy (girlstudy):

I guess?

OpenStudy (mhchen):

Okay, cause we're going to factor in gravity. Hold on give me a sec. to thikn about it.

OpenStudy (mhchen):

 Okay so this is what I'm thinking of. And we have to convert 90miles per hour into m/s

OpenStudy (mhchen):

1,609m in a mile, so 144,810 meters / hour. 3600 seconds in an hour so 40.225 m/s So Fracket = 40.225m/s * 1.5kg = 60.3375N

OpenStudy (mhchen):

Then we use pythagorean theorem to do \[\sqrt{60.3375^2+15^2}\] and that should be it based on my guesses/calculations.

OpenStudy (girlstudy):

Why do we use pythagorean theorem?

OpenStudy (girlstudy):

How did you get 15N in the beggining? where did the 10m/s come from? @mhchen

OpenStudy (girlstudy):

@mathmale

OpenStudy (mhchen):

Um... Force of gravity = mass(kg) * 10m/s^2 (accerlation of gravity)

OpenStudy (mhchen):

The vector components of then x and y directions make part of a right-triangle.

OpenStudy (blazeryder):

isn't gravity 9.81 m/s^2? I guess if you wanted to you could round up to 10, but a Physics teacher might not like that.....

OpenStudy (irishboy123):

 @mhchen ?!?!